Solve each system by addition, if possible.
4x + 6y = 5
8x - 9y = 3

Answer by TimothyLamb(4379)

4x + 6y = 5
8x - 9y = 3
---
-8x - 12y = -10
8x - 9y = 3
---
-21y = -7
---
y = 1/3
4x + 6(1/3) = 5
4x + 2 = 5
4x = 3
x = 3/4
Answer:
x = 3/4
y = 1/3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php